OpenStudy (anonymous):

Ok, I'm lost here! Here are the questions: A drawer contains 4 red socks, 3 white socks, and 3 blue socks. Without looking, you select a sock at random, replace it, and select a second sock at random. What is the probability that the first sock is blue and the second sock is red? A lunch menu consists of 5 different sandwiches, 2 different soups, and 5 different drinks. How many choices are there for ordering a sandwich, a bowl of soup, and a drink? The Burger Diner offers burgers with or without any or all of the following: catsup, lettuce, and mayonnaise. How many different burgers can you order? Math is NOT my strong suit but I would loovvee it if you could explain how to do it! Thank ya!

OpenStudy (anonymous):

sorry I accidentally pressed enter haha!

OpenStudy (anas.p):

Probability of blue socks is 3/10 (3 blue socks out of 10 socks) Probability of red socks is 4/10 (4 red socks out of 10 socks) the question says and therefor the probabilities will be multipled 3/10*4/10 12/100=.12 or 3/25 Because there is selection, we will use combination For ordering a sandwich 5 options are there and you need one so.. 5C1=5 For ordering a soup 2 options are there and you need one so.. 2C1=2 For ordering a drink 5 options are there and you need one so.. 5C1=5 You will multiply all these. 5C1*2C1*5C1=50 This is also about Combination, You can skip all of them 4C0=1 You can have one of them 4C1=4 You can have two of them 4C2=6 You can have three of them 4C3=4 You can have all of them 4C4=1 Multiply all these 1*4*6*4*1=96

OpenStudy (anonymous):

Wow thanks!!! I got it!! :D

OpenStudy (anas.p):

no problem